Indicate the concentration of each ion present in the solution formed by mixing the following.

44.0 mL of 0.110 M Na2SO4 and 25.0 mL of 0.200 M KCl. Assume that the volumes are additive.

Enter your answers, separated by commas, using three significant figures.
Answer #1

In solution :

Na2SO4 (aq) ====> 2 Na+ (aq) + SO42- (aq)

KCl (aq) ====> K+ (aq) + Cl- (aq)

Total Volume of solution after mixing = 44.0 mL + 25.0 mL = 69.0 ml

Conc. of Na+ ion = moles Na+ ion / total volume

= (44 ml* 2 *0.110 M )/ 69.0 ml = 0.140 M

Conc. of K+ ion = moles K+ ion / total volume

= (25 ml *0.200 M )/ 69.0 ml = 0.073 M

Conc. of Cl- ion = moles Cl-ion / total volume

= (25 ml *0.200 M )/ 69.0 ml = 0.073 M

Conc. of SO42- ion = moles SO42- ion / total volume

= (44 ml* 0.110 M )/ 69.0 ml = 0.070 M